# Fixed-Point RNNs: Interpolating from Diagonal to Dense

### v2

@@ Conclusion
B.2 Optimizing Fixed-Point RNNs
@@
One advantage of converging to a fixed-point over general layer looping lies in model training. Since the gradient with respect to $h^{0}$ is not needed, implicit differentiation can be used to avoid storing and backpropagating through the computational graph of the fixed-point iteration, as discussed by Liao et al. (2018), Bai et al. (2019). To see this, consider the Jacobian across $\ell$ iterations $\mathbf{J}_{x}^{\ell}=\frac{\partial f_{\theta}}{\partial x}(x,h^{\ell-1})$. Since $h^{\ell-1}$ depends on $x$ as well, we can recursively express $\mathbf{J}_{x}^{\ell}$ in terms of $\mathbf{J}_{x}^{\ell-1}$ and the Jacobians of a single iteration $\mathbf{J}_{x}(h)=\frac{\partial f_{\theta}}{\partial x}(x,h)$ and $\mathbf{J}_{h}=\frac{\partial f_{\theta}}{\partial h}(x,h)$ by applying the chain rule
$\mathbf{J}_{x}^{\ell}=\mathbf{J}_{x}(h^{\ell-1})+\mathbf{J}_{h^{\ell-1}}\cdot\mathbf{J}_{x}^{\ell-1}.$ (15)
@@
Instead of unrolling, we can implicitly differentiate $h^{*}=f_{\theta}(x,h^{*})$ w.r.t. $x$, which yields $\mathbf{J}_{x}^{*}=\mathbf{J}_{x}(h^{*})+\mathbf{J}_{h^{*}}\cdot\mathbf{J}_{x}^{*}$. Given the conditions on the Lipschitz constant of $f_{\theta}(x,h)$ in $h$, we can assume $\mathbf{J}_{h^{\ell}}$ to be contractive and therefore $\left(\mathbf{I}-\mathbf{J}_{h^{\ell}}\right)$ to be positive definite and invertible. This allows to reformulate as
$\mathbf{J}_{x}^{*}=(\mathbf{I}-\mathbf{J}_{h^{*}})^{-1}\cdot\mathbf{J}_{x}(h^{*}).$ (16)
@@
The case for $\mathbf{J}_{\theta}^{*}$ works analogously. This means that the gradient w.r.t. the input $x$ and parameters $\theta$ can be computed at the fixed-point with the cost of solving $(\mathbf{I}-\mathbf{J}_{h^{*}})^{-1}$. Bai et al. (2021) and Schöne et al. (2025) approximate this inverse using the first terms of the Neumann series, which leads to a truncated backpropagation formulation or phantom gradients, incurring sequential overhead. For iteration with hidden state dependence, we can avoid this inversion altogether with the following workaround:
Theorem B.1.
Let $f_{\theta}(x,h)$ have Lipschitz constant $<1$ and fixed-point $h^{*}$. If the Jacobians $\frac{\partial f_{\theta}}{\partial x}(x,h)$ and $\frac{\partial f_{\theta}}{\partial h}(x,h)$ are equal, then the gradient $\nabla_{\theta}\mathcal{L}\big{(}f_{\theta}(x,h),y\big{)}$ of the loss $\mathcal{L}(\cdot,y)$ for a target $y$ and evaluated at the fixed point $h=h^{*}$ will be a descent direction of $\mathcal{L}\big{(}F_{\theta}(x),y\big{)}$. Proof in App. B.3.
@@
In simple terms, Thm. B.1 shows that parameterizing $f_{\theta}(x,h)$ such that $\mathbf{J}_{x}(h)=\mathbf{J}_{h}$ guarantees optimization progress even if the gradient is computed only at the fixed-point. In practice, we observe that adhering to this condition in the form of hidden state dependence speeds-up the convergence of the model during training.
B.3 Proof for Theorem B.1
@@
We start the proof by setting $\delta:=\frac{\partial\mathcal{L}}{\partial f}$ and $\mathbf{J}_{x}:=\mathbf{J}_{x}(h^{*})$. Then, we can write the backward propagation as $\frac{\partial\mathcal{L}}{\partial x}=\left(\mathbf{J}_{x}^{*}\right)^{\top}\delta$. In order to prove that the gradient computed at the fixed-point is a descent direction, we need to show that $\mathbf{J}_{x}^{\top}\delta$ is in the direction of $\left(\mathbf{J}_{x}^{*}\right)^{\top}\delta$, or in other words, we have $\delta^{\top}\mathbf{J}_{x}^{*}\mathbf{J}_{x}^{\top}\delta\geq 0$. This is equivalent to showing that the symmetric part of the matrix $\mathbf{J}_{x}^{*}\mathbf{J}_{x}^{\top}$ is positive semi-definite.
@@
Now note that from Eq. 16 we have: $\mathbf{J}_{x}^{*}\mathbf{J}_{x}^{\top}=\left(\mathbf{I}-\mathbf{J}_{h}\right)^{-1}\mathbf{J}_{x}\mathbf{J}_{x}^{\top}$. From our assumption $\mathbf{J}_{x}=\mathbf{J}_{h}:=\mathbf{J}$, we need to show that the symmetric part of the matrix $\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}\mathbf{J}^{\top}$ is positive semi-definite. Note that using the Neumann series, we can write:
$\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}=\sum_{i=1}^{\infty}\mathbf{J}^{i}=\mathbf{J}\sum_{i=0}^{\infty}\mathbf{J}^{i}=\mathbf{J}\left(\mathbf{I}-\mathbf{J}\right)^{-1}.$
@@
Therefore, we have $\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}\mathbf{J}^{\top}=\mathbf{J}\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}^{\top}$. Going back to the definition of positive semi-definiteness, we need to show that $\delta^{\top}\mathbf{J}\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\delta>0$ for all $\delta$. Setting $\omega=\mathbf{J}^{\top}\delta$, this is equivalent to having $\omega^{\top}\left(\mathbf{I}-\mathbf{J}\right)^{-1}\omega$. Note that from our assumption for the Lipschitz constant of the function, we have $\left\|\mathbf{J}\right\|_{2}<1$, which means $\left(\mathbf{I}-\mathbf{J}\right)$ has strictly positive eigenvalues. This completes our proof.∎
B.4 Heuristics to reduce the number of fixed-point iterations
The second heuristic involves using a momentum-like update rule to accelerate the convergence of fixed-point iterations for certain sequences. Specifically, we observe that by setting the fixed-point update rule to $h^{\ell+1}=\delta\cdot f_{\theta}(x,h^{\ell})+(1-\delta)\cdot h^{\ell}$ for some $\delta\in[0,1]$, we can accelerate the convergence for certain sequences that are particularly slow to converge. Since this update rule can result in a biased approximation of the fixed-point, we implement a patience-based system that starts with $\delta=1$, and reduces the value of $\delta$ exponentially when the residues fail to improve.
@@ C.1 Mamba: Selective SSMs
$\mathbf{h}_{t}=\boldsymbol{\lambda}_{t}\odot\mathbf{h}_{t-1}+\mathbf{B}_{t}\left(\Delta_{t}x_{t}\right)^{\top},$ (17)
@@
where $\boldsymbol{\lambda}_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$ is an input-dependent state transition vector, $\mathbf{B}_{t}\in\mathbb{R}^{d_{\text{state}}}$ an input transition vector, $x_{t}\in\mathbb{R}^{d_{\text{inner}}}$ the input, and $\Delta_{t}\in\mathbb{R}^{d_{\text{inner}}\times d_{\text{inner}}}$ a diagonal matrix which acts an input normalization term. The matrices are parameterized as:
$\boldsymbol{\lambda}_{t}$ $=\exp\left(-\boldsymbol{\lambda}_{\log}\Delta_{t}\right),$ $\boldsymbol{\lambda}_{\log}$ $=\exp\left(\omega\right),$
$\Delta_{t}$ $=\text{diag}\left(\text{softplus}\left(\mathbf{W}_{\Delta}x_{t}+b_{\Delta}\right)\right),$ $\mathbf{B}_{t}$ $=\mathbf{W}_{\mathbf{B}}x_{t},$
@@
with $\omega\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$, $\mathbf{W}_{\Delta}\in\mathbb{R}^{d_{\text{inner}}\times d_{\text{inner}}}$, $\mathbf{W}_{\mathbf{B}}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$, and $b_{\Delta}\in\mathbb{R}^{d_{\text{inner}}}$. The output of a Mamba block $y_{t}\in\mathbb{R}^{d_{\text{inner}}}$ is a contraction of the matrix hidden state with $\mathbf{C}_{t}\in\mathbb{R}^{d_{\text{state}}}$
$y_{t}^{\top}=\mathbf{C}_{t}^{\top}\mathbf{h}_{t},\quad\mathbf{C}_{t}=\mathbf{W}_{\mathbf{C}}x_{t},$
@@
for $\mathbf{W}_{\mathbf{C}}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$. Note that Mamba proposes a skip connection of $y_{t}+\mathbf{D}\odot x_{t}$, where $\mathbf{D}\in\mathbb{R}^{d_{\text{inner}}}$ is an input-independent vector. Finally, the model output is usually scaled by a gated linear unit (GLU) as $\tilde{y}_{t}=\mathbf{G}_{t}\odot y_{t}$, where $\mathbf{G}_{t}=\text{SiLU}\left(\mathbf{W}_{G}x_{t}\right)$ is a non-linear function of the input.
C.2 FP-Parametrization

### v3

@@ Conclusion
B.2 Effect of normalization factor $(\mathbf{I}-\mathbf{\Lambda}_{t})$ on class of matrices $\mathbf{A}_{t}$
For the sake of exposition, the introduction of the implicit parametrization in Sec. 3.1 did not consider input normalization $(\mathbf{I}-\mathbf{\Lambda}_{t})$. However, as discussed in Sec. 3.2 this is a crucial component to stabilize the recurrence in time. To derive the representable dense matrices $\mathbf{A}_{t}$ in the presence of the normalization factor $(\mathbf{I}-\mathbf{\Lambda}_{t})$, let us start by assuming a fixed-point was found according to Thm. 3.2:
$\mathbf{h}^{*}_{t}$ $=\mathbf{\Lambda}_{t}\mathbf{h}^{*}_{t-1}+(\mathbf{I}-\mathbf{\Lambda}_{t})(\mathbf{Q}_{t}\mathbf{B}_{t}\mathbf{x}_{t}+(\mathbf{I}-\mathbf{Q}_{t})\mathbf{h}_{t}^{*}$
$=\mathbf{\Lambda}_{t}\mathbf{h}^{*}_{t-1}+(\mathbf{I}-\mathbf{\Lambda}_{t})\mathbf{Q}_{t}\mathbf{B}_{t}\mathbf{x}_{t}+(\mathbf{I}-\mathbf{\Lambda}_{t})(\mathbf{I}-\mathbf{Q}_{t})\mathbf{h}_{t}^{*}.$
Rearranging the terms allows to move $\mathbf{h}^{*}_{t}$ to the other side
$(\mathbf{I}-(\mathbf{I}-\mathbf{\Lambda}_{t})(\mathbf{I}-\mathbf{Q}_{t}))\mathbf{h}^{*}_{t}=\mathbf{\Lambda}_{t}\mathbf{h}^{*}_{t-1}+(\mathbf{I}-\mathbf{\Lambda}_{t})\mathbf{Q}_{t}\mathbf{B}_{t}\mathbf{x}_{t}.$
Moving $(\mathbf{I}-(\mathbf{I}-\mathbf{\Lambda}_{t})(\mathbf{I}-\mathbf{Q}_{t}))$ back to other side yields
$\mathbf{A}_{t}$ $=(\mathbf{I}-(\mathbf{I}-\mathbf{\Lambda}_{t})(\mathbf{I}-\mathbf{Q}_{t}))^{-1}\mathbf{\Lambda}_{t}$
$=\left(\mathbf{\Lambda}_{t}^{-1}(\mathbf{I}-(\mathbf{I}-\mathbf{\Lambda}_{t})(\mathbf{I}-\mathbf{Q}_{t}))\right)^{-1}$
$=\left(\mathbf{\Lambda}_{t}^{-1}-(\mathbf{\Lambda}_{t}^{-1}-\mathbf{I})(\mathbf{I}-\mathbf{Q}_{t})\right)^{-1}$
$=\left(\mathbf{I}+(\mathbf{\Lambda}_{t}^{-1}-\mathbf{I})\mathbf{Q}_{t}\right)^{-1}$
Following the standard assumptions that $\mathbf{0}\preceq\mathbf{\Lambda}_{t},\mathbf{Q}_{t}\preceq\mathbf{I}$, the matrix $(\mathbf{I}+(\mathbf{\Lambda}_{t}^{-1}-\mathbf{I})\mathbf{Q}_{t})$ is full rank and $\succeq\mathbf{I}$. Therefore its inverse $\mathbf{A}_{t}$ exists and is contractive. The expressivity of $\mathbf{A}_{t}$ is only limited if $\mathbf{\Lambda}_{t}\approx\mathbf{I}$. This however would also be problematic for diagonal SSM and therefore the Mamba initialization is bias towards $\mathbf{\Lambda}_{t}\prec\mathbf{I}$. Thus, the normalization does not pose a significant problem for the expressivity of $\mathbf{A}_{t}$ in practice.
B.3 Implicit Differentiation for Optimizing Fixed-Point RNNs
@@
One advantage of converging to a fixed-point over general layer looping lies in model training. Since the gradient with respect to $\mathbf{h}^{0}$ is not needed, implicit differentiation can be used to avoid storing and backpropagating through the computational graph of the fixed-point iteration, as discussed by Liao et al. (2018), Bai et al. (2019). To see this, consider the Jacobian across $\ell$ iterations $\mathbf{J}_{\mathbf{x}}^{\ell}=\frac{\partial f_{\boldsymbol{\theta}}}{\partial\mathbf{x}}(\mathbf{x},\mathbf{h}^{\ell-1})$. Since $\mathbf{h}^{\ell-1}$ depends on $\mathbf{x}$ as well, we can recursively express $\mathbf{J}_{\mathbf{x}}^{\ell}$ in terms of $\mathbf{J}_{\mathbf{x}}^{\ell-1}$ and the Jacobians of a single iteration $\mathbf{J}_{\mathbf{x}}(\mathbf{h})=\frac{\partial f_{\boldsymbol{\theta}}}{\partial\mathbf{x}}(\mathbf{x},\mathbf{h})$ and $\mathbf{J}_{\mathbf{h}}=\frac{\partial f_{\boldsymbol{\theta}}}{\partial\mathbf{h}}(\mathbf{x},\mathbf{h})$ by applying the chain rule
$\mathbf{J}_{\mathbf{x}}^{\ell}=\mathbf{J}_{\mathbf{x}}(\mathbf{h}^{\ell-1})+\mathbf{J}_{\mathbf{h}^{\ell-1}}\cdot\mathbf{J}_{\mathbf{x}}^{\ell-1}.$ (16)
@@
Instead of unrolling, we can implicitly differentiate $\mathbf{h}^{*}=f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h}^{*})$ w.r.t. $\mathbf{x}$, which yields $\mathbf{J}_{\mathbf{x}}^{*}=\mathbf{J}_{\mathbf{x}}(\mathbf{h}^{*})+\mathbf{J}_{\mathbf{h}^{*}}\cdot\mathbf{J}_{\mathbf{x}}^{*}$. Given the conditions on the Lipschitz constant of $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ in $\mathbf{h}$, we can assume $\mathbf{J}_{\mathbf{h}^{\ell}}$ to be contractive and therefore $\left(\mathbf{I}-\mathbf{J}_{\mathbf{h}^{\ell}}\right)$ to be positive definite and invertible. This allows to reformulate as
$\mathbf{J}_{\mathbf{x}}^{*}=(\mathbf{I}-\mathbf{J}_{\mathbf{h}^{*}})^{-1}\cdot\mathbf{J}_{\mathbf{x}}(\mathbf{h}^{*}).$ (17)
@@
The case for $\mathbf{J}_{\boldsymbol{\theta}}^{*}$ works analogously. This means that the gradient w.r.t. the input $x$ and parameters $\boldsymbol{\theta}$} can be computed at the fixed-point with the cost of solving $(\mathbf{I}-\mathbf{J}_{\mathbf{h}^{*}})^{-1}$. Bai et al. (2021) and Schöne et al. (2025) approximate this inverse using the first terms of the Neumann series, which leads to a truncated backpropagation formulation or phantom gradients, incurring sequential overhead. For iteration with hidden state dependence, we can avoid this inversion altogether with Thm. 3.2:
See 3.2
@@
In simple terms, Thm. 3.2 shows that parameterizing $f_{\boldsymbol{\theta}}(\mathbf{x},\mathbf{h})$ such that $\mathbf{J}_{\mathbf{x}}(\mathbf{h})=\mathbf{J}_{\mathbf{h}}$ guarantees optimization progress even if the gradient is computed only at the fixed-point. In practice, we observe that adhering to this condition in the form of hidden state dependence speeds-up the convergence of the model during training.
B.4 Proof for Theorem 3.2 (Gradient of $f_{\theta}(\mathbf{x},\mathbf{h})$ is a descent direction of $F_{\theta}(\mathbf{x})$)
@@
We start the proof by setting $\boldsymbol{\delta}:=\frac{\partial\mathcal{L}}{\partial f}$ and $\mathbf{J}_{\mathbf{x}}:=\mathbf{J}_{\mathbf{x}}(\mathbf{h}^{*})$. Then, we can write the backward propagation as $\frac{\partial\mathcal{L}}{\partial\mathbf{x}}=\left(\mathbf{J}_{\mathbf{x}}^{*}\right)^{\top}\boldsymbol{\delta}$. In order to prove that the gradient computed at the fixed-point is a descent direction, we need to show that $\mathbf{J}_{\mathbf{x}}^{\top}\boldsymbol{\delta}$ is in the direction of $\left(\mathbf{J}_{\mathbf{x}}^{*}\right)^{\top}\boldsymbol{\delta}$, or in other words, we have $\boldsymbol{\delta}^{\top}\mathbf{J}_{\mathbf{x}}^{*}\mathbf{J}_{\mathbf{x}}^{\top}\boldsymbol{\delta}\geq 0$. This is equivalent to showing that the symmetric part of the matrix $\mathbf{J}_{\mathbf{x}}^{*}\mathbf{J}_{\mathbf{x}}^{\top}$ is positive semi-definite.
@@
Now note that from Eq. 17 we have: $\mathbf{J}_{\mathbf{x}}^{*}\mathbf{J}_{\mathbf{x}}^{\top}=\left(\mathbf{I}-\mathbf{J}_{\mathbf{h}}\right)^{-1}\mathbf{J}_{\mathbf{x}}\mathbf{J}_{\mathbf{x}}^{\top}$. From our assumption $\mathbf{J}_{\mathbf{x}}=\mathbf{J}_{\mathbf{h}}:=\mathbf{J}$, we need to show that the symmetric part of the matrix $\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}\mathbf{J}^{\top}$ is positive semi-definite. Note that $(\mathbf{I}-\mathbf{J})^{-1}$ and $\mathbf{J}$ commute by application of the Neumann series
$\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}=\sum_{i=1}^{\infty}\mathbf{J}^{i}=\mathbf{J}\sum_{i=0}^{\infty}\mathbf{J}^{i}=\mathbf{J}\left(\mathbf{I}-\mathbf{J}\right)^{-1},$
@@
which yields $\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}\mathbf{J}^{\top}=\mathbf{J}\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}^{\top}$. Going back to the definition of positive semi-definiteness, we need to show that $\boldsymbol{\delta}^{\top}\mathbf{J}\left(\mathbf{I}-\mathbf{J}\right)^{-1}\mathbf{J}^{\top}\boldsymbol{\delta}>0$ for all $\boldsymbol{\delta}$. Setting $\boldsymbol{\omega}=\mathbf{J}^{\top}\boldsymbol{\delta}$, this is equivalent to having $\boldsymbol{\omega}^{\top}\left(\mathbf{I}-\mathbf{J}\right)^{-1}\boldsymbol{\omega}$. Note that from our assumption for the Lipschitz constant of the function, we have $\left\|\mathbf{J}\right\|_{2}<1$, which means $\left(\mathbf{I}-\mathbf{J}\right)$ and $\left(\mathbf{I}-\mathbf{J}\right)^{-1}$ have strictly positive eigenvalues. This completes our proof.∎
@@ C.1 Mamba: Selective SSMs
$\mathbf{H}_{t}=\boldsymbol{\lambda}_{t}\odot\mathbf{H}_{t-1}+\mathbf{b}_{t}\left(\Delta_{t}\mathbf{x}_{t}\right)^{\top},$ (18)
@@
where $\boldsymbol{\lambda}_{t}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$ is an input-dependent state transition vector, $\mathbf{b}_{t}\in\mathbb{R}^{d_{\text{state}}}$ an input transition vector, $\mathbf{x}_{t}\in\mathbb{R}^{d_{\text{inner}}}$ the input, and $\Delta_{t}\in\mathbb{R}^{d_{\text{inner}}\times d_{\text{inner}}}$ a diagonal matrix which acts an input normalization term. The matrices are parameterized as:
$\boldsymbol{\lambda}_{t}$ $=\exp\left(-\boldsymbol{\lambda}_{\log}\Delta_{t}\right),$ $\boldsymbol{\lambda}_{\log}$ $=\exp\left(\boldsymbol{\omega}\right),$
$\Delta_{t}$ $=\text{diag}\left(\text{softplus}\left(\mathbf{W}_{\Delta}\mathbf{x}_{t}+b_{\Delta}\right)\right),$ $\mathbf{b}_{t}$ $=\mathbf{W}_{\mathbf{b}}\mathbf{x}_{t},$
@@
with $\boldsymbol{\omega}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$, $\mathbf{W}_{\Delta}\in\mathbb{R}^{d_{\text{inner}}\times d_{\text{inner}}}$, $\mathbf{W}_{\mathbf{b}}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$, and $b_{\Delta}\in\mathbb{R}^{d_{\text{inner}}}$. The output of a Mamba block $\mathbf{y}_{t}\in\mathbb{R}^{d_{\text{inner}}}$ is a contraction of the matrix hidden state with $\mathbf{c}_{t}\in\mathbb{R}^{d_{\text{state}}}$
$\mathbf{y}_{t}^{\top}=\mathbf{c}_{t}^{\top}\mathbf{H}_{t},\quad\mathbf{c}_{t}=\mathbf{W}_{\mathbf{c}}\mathbf{x}_{t},$
@@
for $\mathbf{W}_{\mathbf{c}}\in\mathbb{R}^{d_{\text{state}}\times d_{\text{inner}}}$. Note that Mamba proposes a skip connection of $\mathbf{y}_{t}+\mathbf{D}\odot\mathbf{x}_{t}$, where $\mathbf{D}\in\mathbb{R}^{d_{\text{inner}}}$ is an input-independent vector. Finally, the model output is usually scaled by a gated linear unit (GLU) as $\mathbf{\tilde{y}}_{t}=\mathbf{g}_{t}\odot\mathbf{y}_{t}$, where $\mathbf{g}_{t}=\text{SiLU}\left(\mathbf{W}_{\mathbf{g}}\mathbf{x}_{t}\right)$ is a non-linear function of the input.
C.2 FP-Mamba Parametrization